OpenStudy (anonymous):

eight times the reciprocal of a # equals 2 times the reciprocal of 9. find the number

Directrix (directrix):

@lashell123 Let x be the #. Then, the reciprocal of the # is 1/x where x is not 0. eight times the reciprocal of a # equals 2 times the reciprocal of 9. find the number ------------------ 8*(1/x) = 2*(1/9) 8/x = 2/9 Solve for x.